OpenStudy (anonymous):

-3x + 2y = 8 y = -3x/2 + 4

OpenStudy (anonymous):

aha sike i mean 1. Choose one of the groups of equations from the table below. Indicate which group you choose in your word processing file. Group 1 Group 2 Group 3 Group 4 y = x + 4 –3x + 2y = 8 y = x – 2 –2x + 3y = -6 y = x + 3 –x + 3y = 9 y = x + 5 –x + 6y = 30

OpenStudy (anonymous):

2. Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. 3. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 4. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. 5. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 6. In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences omg i need help im behind :/